Manufacturers Southern leased high-tech electronic equipment from Edison Leasing on January 1, 2018. Edison purchased the equipment from International Machines at a cost of $123,.288. (FV of $1. PV of $1. FVA of S1. PVA of S1. FVAD of $1 and PVAD of S) (Use appropriate factorls) from the tables provided.) Related Information: Lease term Quarterly rental payments Economic life of asset Fair value of asset Implicit interest rate (Also lessees increnentol borrowing rate) 2 years (8 quarterly periods) $16,500 at the beginning of each period 2 years $123,288 8X Required: Prepare a lease amortization schedule and appropriate entries for Manufacturers Southern from the beginning of the lease through January 1,2019 Depreciation is recorded at the end of each fiscal year (December 31) on a straight-ine basis. Complete this question by entering your answers in the tabs below. Amort General Journal Schedule Prepare a lease amortization schedule for the term of the lease for Manufacturers Southern from the beginning of the lease through January 1, 2019. Depreciation is recorded at the end of each fiscal year (December 31) on a straight-line basis. (Enter your answers in whole dollars and not in milltions. Round your intermedilate and final answers to nearest whole dollar Enter all amounts as positive values.) Decrease in Lease Balance ayment batePayments InterestBalance
Prepare a lease amortization schedule and appropriate entries for Manufacturers Southern from the beginning of the lease through January 1, 2019. Depreciation is recorded at the end of each fiscal year (December 31) on a straight-line basis Complete this question by entering your answers in the tabs below. 10 points General Journal Amort Book Hint Print References Prepare a lease amortization schedule for the term of the lease for Manufacturers Southern from the beginning of the lease through January 1, 2019. Depreclation is recorded at the end of each fiscal year (December 31) on a straight-line basis Enter your answers in whole dollars and not in millions. Round your intermediate and final answers to nearest whole dollar Enter all amounts as positive values.) E fective Decrease in Lease Balance Payment Date 1/012018 04/01/2018 71012018 10/01/2018 01/01/2019 04/01/2019 07/01/2019 10/01201 Total General Journal>
Answer #1
Payment Date Lease Payment Interest Rate Previous quarter balance Effective Interest Decrease in Balance Lease Balance
01-01-2018 a b c d=b*c e=a-d 123,288.00
01-01-2018 16,500.00 16,500.00 106,788.00
04-01-2018 16,500.00 2% 106,788.00 2135.76 14,364.24 92,423.76
07-01-2018 16,500.00 2% 92,423.76 1848.48 14,651.52 77,772.24
10-01-2018 16,500.00 2% 77,772.24 1555.44 14,944.56 62,827.68
01-01-2019 16,500.00 2% 62,827.68 1256.55 15,243.45 47,584.23
04-01-2019 16,500.00 2% 47,584.23 951.68 15,548.32 32,035.92
07-01-2019 16,500.00 2% 32,035.92 640.72 15,859.28 16,176.64
10-01-2019 16,500.00 2% 16,176.64 323.53 16,176.47 0.17
Total 132,000.00 8,712.17 123,287.83